# A data augmentation strategy for deep neural networks with application to epidemic modelling

### v1

@@ Abstract
In this work, we integrate the predictive capabilities of compartmental disease dynamics models with machine learning ability to analyze complex, high-dimensional data and uncover patterns that conventional models may overlook. Specifically, we present a proof of concept demonstrating the application of data-driven methods and deep neural networks to a recently introduced SIR-type model with social features, including a saturated incidence rate, to improve epidemic prediction and forecasting. Our results show that a robust data augmentation strategy trough suitable data-driven models can improve the reliability of Feed-Forward Neural Networks (FNNs) and Nonlinear Autoregressive Networks (NARs), making them viable alternatives to Physics-Informed Neural Networks (PINNs). This approach enhances the ability to handle nonlinear dynamics and offers scalable, data-driven solutions for epidemic forecasting, prioritizing predictive accuracy over the constraints of physics-based models. Numerical simulations of the post-lockdown phase of the COVID-19 epidemic in Italy and Spain validate our methodology.
@@ 1 Introduction
The COVID-19 pandemic has underscored the need for accurate and computationally efficient predictive models that not only describe the spread of infectious diseases but also provide fast and reliable forecasts to guide public health interventions. A key challenge in epidemic modeling is balancing mechanistic realism with computational tractability. Among the most widely used approaches are compartmental models, such as the classical Susceptible-Infectious-Recovered (SIR) models, which describe disease dynamics by partitioning the population into epidemiologically relevant compartments [16, 29, 32]. Specifically, these models categorize individuals as susceptible (capable of contracting the disease), infected (currently infected and capable of transmitting the disease), and recovered (immune or isolated from the population). Despite their simplicity and effectiveness in capturing some fundamental epidemic trends, classical SIR models assume constant transmission and recovery rates, limiting their ability to represent the complex, nonlinear, and time-varying nature of real-world epidemics.
@@
Feed-Forward Neural Networks (FNNs) and Nonlinear Autoregressive Networks (NARs) also offer promising approaches for epidemiological predictions [27, 2, 34]. FNNs map inputs directly to outputs through layers of nonlinear transformations, making them effective for tasks such as interpolation and short-term forecasting. This approach has been applied across various domains, including early recolonization detection in Parkinson disease, biological parameter estimation in disease models, COVID-19 data analysis, musculoskeletal disorder risk assessment, and statistical modeling for selecting optimal FNN architectures [24, 44, 38, 42, 17, 36]. However, FNNs do not account for temporal dependencies, limiting their ability to model the evolution of epidemics over time.
@@
In this work, we introduce a hybrid epidemic forecasting approach that integrates mechanistic epidemiological models with data-driven FNNs and NARs, creating a surrogate model capable of generalizing across different epidemic scenarios. This approach enables fast, data-driven epidemic forecasting while preserving the interpretability of traditional mechanistic models. The resulting methodology is applied to analyze the post-lockdown phase of the COVID-19 epidemic in Italy and Spain, leveraging real-world data and enhanced computational strategies to improve predictive accuracy. This framework enables us to augment real-world epidemic data with synthetic simulations generated from a generalized SIR model, which includes a saturated incidence rate to better reflect intervention effects and behavioral changes. The main idea consists in addressing the limitations posed by sparse real-world data by generating additional synthetic information using a generalized SIR data driven model. Specifically, we focus on an extension known as the social-SIR model with a saturated incidence rate, referred to here as the generalized SIR model in the following. This model modifies the transmission dynamics [23] by allowing the infection rate to vary with time and the number of infected individuals. Such modifications enable the incorporation of external interventions, such as lockdowns or vaccination campaigns, as well as psychological and behavioral adaptations [14].
@@
The remainder of this paper is structured as follows. In Section 2, we recall some standard features about the classical SIR model and we introduce the social-SIR model with a saturated incidence rate used to produce synthetic data. Section 3 defines the optimization problems used to estimate parameters and the incidence function during the pre- and post-lockdown phases in Italy and Spain. In Section 4, we present the neural network structures, distinguishing between FNNs and NARs and we perform some detailed numerical experiments to validate our strategy. Finally, Section 5 summarizes our findings and outlines potential future research directions.
@@ 2 Problem setting
The governing equations of the standard SIR model, which describes the spread of an infectious disease in a population of $N$ individuals divided into three compartments, Susceptible ($S$), Infected ($I$), and Recovered ($R$), are given by
@@
The above model despite of its simplicity it has been shown to be able to capture some features related to the spread of an epidemic. However, as expected, it also suffers of some drawbacks due to the inability to describe more complex structured solutions which has been observed for instance during the last COVID-19 epidemic [26]. In particular, recent studies have highlighted the crucial role of transmission coefficients in relation to social behavior, which significantly influences pathogen spread [5]. In this research direction, the works in [5, 23] explored how controlling the number of daily social contacts can guide partial lockdown strategies. This is type of phenomena we address in this work and which cannot be captured with the system (2.1). Building on this perspective, some of the authors of the present work have developed in the recent past a new mathematical framework linking social contact distributions with disease transmission in multi-agent systems in [23]. This is achieved by integrating a classical compartmental epidemiological model with kinetic equations governing social contact formation. The resulting model is given by
@@
Alternatively, setting $m=1,k=2$ results in the MonodÄìHaldane (MÄìH) function,
@@
The so-called social-SIR model (2.2) exhibits a similar incidence rate behavior, extending the form in (2.5) to account for the impact of protective measures such as lockdown strategies and infection control. Unlike classical models, this generalized framework incorporates a time-dependent, saturated incidence rate to capture the cumulative effects of infection dynamics and behavioral adaptations in response to an epidemic. This model has previously been successfully employed to analyze the effects of partial lockdown strategies in various European countries. Given that we choose in the following, for demonstrating the proposed data augmentation strategy, a dataset corresponding to a period marked by COVID-19 lockdowns, we opted to use this model in our study. It is worth to mention that, the complexity introduced by lockdown measures, makes the forecasting process more challenging compared to standard scenarios. Therefore, in the following, we employ system (2.2) to develop our data-driven strategy and the associated deep learning framework, enabling epidemic forecasting.
@@ 3 On the the construction of a data driven social-SIR model
In the context of epidemic prediction, mathematical models such as those presented in the previous section provide valuable insights into the spread of infectious diseases. However, these models often rely on predefined assumptions and parameters that may not fully capture the complexities of real-world epidemics. Neural networks offer a powerful alternative by learning directly from data, enabling them to model complex, nonlinear relationships that traditional epidemic models may struggle to represent. However, one major limitation of training neural networks exclusively on available data is that, while they can effectively interpolate within a given time interval, they struggle to make accurate predictions beyond the observed data points [33, 2, 9]. To address this issue, we propose an approach that combines a mathematical epidemiological modelÄîthe social-SIR model (2.2)Äîwith real outbreak data. The key idea is to generate additional synthetic information by solving this generalized SIR model, thereby filling the gaps between available real or measured data points. With this enriched dataset, one can train a neural network more effectively to predict future infection trends in a given population. As we demonstrate later, this approach leverages both real and synthetic data, significantly improving predictive accuracy and enhancing the modelÄôs ability to forecast epidemic dynamics. However, to implement this strategy, the model must be capable of interpolating through the experimental data points. Therefore, in the following section, we discuss how the parameters in (2.2) are defined.
@@ 3.1 Parameters estimation
To develop our approach, we focus on the COVID-19 outbreak in two European countries, Italy and Spain, during the 2020 pandemic. Given that both countries implemented restrictive measures during this period, we proceed as follows. We define two optimization phases: in the first phase, we estimate the parameters $\beta$ and $\gamma$ for the pre-lockdown period using the model (2.2) with $H(t,I(t))=1$; in the second phase, we estimate the function $H(t,I(t))$ from post-lockdown data.
@@
In Figure 1 on the left the evolution of the dynamics in the pre-lockdown phase in Italy obtained by solving the social-SIR model (2.2) with $H(t,I(t))=1$, using the estimated parameters $\beta=\beta_{est}^{It}$ and $\gamma=\gamma_{est}^{It}$ in (3.4). This is compared, in the same image, to the experimental data from the pre-lockdown phase, spanning up to time $t_{l}=14$ with a time step $\Delta t=1$ corresponding to $14$ days with $1$ day interval between the measurements. The estimated dynamics aligns quite well with the available data, as confirmed by the embedded plot which shows the pointwise error over time between the solution of the generalized SIR model and the data at disposal, i.e.
@@
Once $\beta$ and $\gamma$ have been estimated, we proceed to determine the pointwise values $H(t,I(t))$ in (2.2) in such a way that the model fits as well as possible the post-lockdown data, i.e., data in the interval $[t_{l},T]$, where $T$ (set to July 1st) is the final observation time for both countries. To this end, we solve the following new optimization problem
@@
where $\mathbb{P}_{[0,1]}(\cdot)$ denotes the projection over the interval $[0,1]$.
The results of this study are reported in Figure 2 where on the left we show the post-lockdown experimental data and the solution of the generalized SIR model with the contact function $H$ as in (3.12). The estimated and true solution are in good agreement as shown in the embedded plot which represents the corresponding pointwise error computed as in (3.6). On the right, we show the estimated values of $H^{n}$ computed as in (3.12) as a function of time. The function $H$ exhibits oscillations over time, suggesting potential measurement errors in the data. First row, Italy. Second row, Spain.
Figure 2: Post-lockdown phase. On the left the dynamics of the infected population computed by solving the generalized SIR model (2.2) with $\beta$ and $\gamma$ as in (3.4) (Italy), and as in (3.5) (Spain) compared with the experimental data. In the embedded plot the pointwise error between the two infected populations computed as in (3.6). On the right the contact function $H$ computed as in (3.12). First row, Italy. Second row, Spain.
@@ 4.1 Neural network for epidemic prediction
To evaluate the performances of FFNs and NARs in epidemic prediction, we conduct some numerical experiments using both real and synthetic data. The real data covers the period from the lockdown start date (in Italy March 9th, in Spain March 14th) to July 1st. These data are available on a daily basis, corresponding to a time step of $\Delta t_{r}=1$. The synthetic data are generated by solving the generalized SIR model given in (2.2), with $\beta$ and $\gamma$ as in (3.4) (Italy), and as in (3.5) (Spain) over the same time interval, using a finer time step $\Delta t_{s}=0.2$, and with $H$ estimated by solving the minimization problem in (3.7). Since $H$ is originally defined at daily intervals as in (3.12), we use linear interpolation to extend its definition to the finer grid. Both networks are trained using the Levenberg-Marquardt optimizer for $500$ epochs. In the case of real data we set $\omega_{r}=1$ and $\omega_{s}=0$, while in the case of synthetic data we set $\omega_{r}=0$ and $\omega_{s}=1$ in (4.3).
@@
We design a FFN with three fully connected layers, each consisting of $10$ nodes, and use the tanh activation function, relying on the use of the Matlab function feedforwardnet. The input to the network is the time interval $[t_{l},T]$, being $t_{l}$ the lockdown starting date and $T$ the final time supposed to be July 1st. The output is the normalized number of infected $I(t)$, which corresponds to either real data or synthetic data generated as described above. We train the neural network using the Matlab function train. To evaluate the neural network’s performance, we first compute its solution over the training set for both Italy (first row) and Spain (second row). Specifically, we consider the time interval $[t_{l},T]$ with a time step $\Delta t_{r}=1$. In the left panel of Figure 3 and of Figure 4, we compare the solution obtained using a neural network trained on both real and synthetic data, with the available data for respectively the Italian and the Spanish case. The neural network provides a good approximation of the solution, making it well-suited for interpolation tasks. Next, to assess its predictive capabilities, we focus on the time interval $(T,T_{test}]$, where $T$ (July 1st) is the starting time and $T_{test}$ (July 15th) is the final time, using the same daily discretization. The right panels of Figures 3 and 4 present a comparison between the network’s predictions for the test set and the daily real data available in that period. Additionally, the figures display the pointwise error between the neural network’s outputs and the real data, calculated as in (3.6). The standard FNN as expected faces challenges in forecasting future dynamics and the data augmentation does not provide a clear improvement in the results.

### v2

@@ Abstract
In this work, we integrate the predictive capabilities of compartmental disease dynamics models with machine learning’s ability to analyze complex, high-dimensional data and uncover patterns that conventional models may overlook. Specifically, we present a proof of concept demonstrating the application of data-driven methods and deep neural networks to a recently introduced Susceptible-Infected-Recovered type model with social features, including a saturated incidence rate, to improve epidemic prediction and forecasting. Our results show that a robust data augmentation strategy trough suitable data-driven models can improve the reliability of Feed-Forward Neural Networks and Nonlinear Autoregressive Networks, providing a complementary strategy to Physics-Informed Neural Networks, particularly in settings where data augmentation from mechanistic models can enhance learning. This approach enhances the ability to handle nonlinear dynamics and offers scalable, data-driven solutions for epidemic forecasting, prioritizing predictive accuracy over the constraints of physics-based models. Numerical simulations of the lockdown and post-lockdown phase of the COVID-19 epidemic in Italy and Spain validate our methodology.
@@ 1 Introduction
The COVID-19 pandemic has underscored the need for accurate and computationally efficient predictive models that not only describe the spread of infectious diseases but also provide fast and reliable forecasts to guide public health interventions. A key challenge in epidemic modeling is balancing mechanistic realism with computational tractability. Among the most widely used approaches are compartmental models, such as the classical Susceptible-Infected-Recovered (SIR) models, which describe disease dynamics by partitioning the population into epidemiologically relevant compartments [16, 29, 32]. Specifically, these models categorize individuals as susceptible (capable of contracting the disease), infected (currently infected and capable of transmitting the disease), and recovered (immune or isolated from the population, i.e. people that no longer contribute to disease transmission). Despite their simplicity and effectiveness in capturing some fundamental epidemic trends, classical SIR models assume constant transmission and recovery rates, limiting their ability to represent the complex, nonlinear, and time-varying nature of real-world epidemics.
@@
Feed-Forward Neural Networks (FNNs) and Nonlinear Autoregressive Networks (NARs) also offer promising approaches for epidemiological predictions [27, 2, 34]. FNNs map inputs directly to outputs through layers of nonlinear transformations, making them effective for tasks such as interpolation and short-term forecasting. This approach has been applied across various domains, including early recolonization detection in Parkinson’s disease, biological parameter estimation in disease models, COVID-19 data analysis, musculoskeletal disorder risk assessment, and statistical modeling for selecting optimal FNN architectures [24, 44, 38, 42, 17, 36]. However, FNNs do not account for temporal dependencies, limiting their ability to model the evolution of epidemics over time.
@@
In this work, we introduce a hybrid epidemic forecasting approach that integrates mechanistic epidemiological models with data-driven FNNs and NARs, creating a surrogate model capable of generalizing across different epidemic scenarios. This approach enables fast, data-driven epidemic forecasting while preserving the interpretability of traditional mechanistic models. The resulting methodology is applied to analyze the lockdown and post-lockdown phase of the COVID-19 epidemic in Italy and Spain, leveraging real-world data and enhanced computational strategies to improve predictive accuracy. This framework enables us to augment real-world epidemic data with synthetic simulations generated from a generalized SIR model, which includes a saturated incidence rate to better reflect intervention effects and behavioral changes. The main idea consists in addressing the limitations posed by sparse real-world data by generating additional synthetic information using a generalized SIR data driven model. Specifically, we focus on an extension known as the social-SIR model with a saturated incidence rate, referred to here as the generalized SIR model in the following. This model modifies the transmission dynamics [23] by allowing the infection rate to vary with time and the number of infected individuals. Such modifications enable the incorporation of external interventions, such as lockdowns or vaccination campaigns, as well as psychological and behavioral adaptations [14].
@@
The remainder of this paper is structured as follows. In Section 2, we recall some standard features about the classical SIR model and we introduce the social-SIR model with a saturated incidence rate used to produce synthetic data. Section 3 defines the optimization problems used to estimate parameters and the incidence function during the pre-lockdown, lockdown and post-lockdown phases in Italy and Spain. In Section 4, we present the neural network structures, distinguishing between FNNs and NARs, and we perform some detailed numerical experiments to validate our strategy. Finally, Section 5 summarizes our findings and outlines potential future research directions.
@@ 2 Problem setting
The governing equations of the standard SIR model, which describes the spread of an infectious disease in a population of $N$ individuals divided into three compartments—Susceptible ($S$), Infected ($I$), and Recovered ($R$)—are given by
@@
Despite its simplicity, the above model has been shown to be able to capture some features related to the spread of an epidemic. However, as expected, it also suffers of some drawbacks due to the inability to describe more complex structured solutions which has been observed for instance during the last COVID-19 epidemic [26]. In particular, recent studies have highlighted the crucial role of transmission coefficients in relation to social behavior, which significantly influences pathogen spread [5]. In this research direction, the works in [5, 23] explored how controlling the number of daily social contacts can guide partial lockdown strategies. This is type of phenomena we address in this work and which cannot be captured with the system (2.1). Building on this perspective, some of the authors of the present work have developed in the recent past a new mathematical framework linking social contact distributions with disease transmission in multi-agent systems in [23]. This is achieved by integrating a classical compartmental epidemiological model with kinetic equations governing social contact formation. The resulting model is given by
@@
Alternatively, setting $m=1,k=2$ results in the Monod–Haldane (M–H) function,
@@
The so-called social-SIR model (2.2) exhibits a similar incidence rate behavior, extending the form in (2.5) to account for the impact of protective measures such as lockdown strategies and infection control. Unlike classical models, this generalized framework incorporates a time-dependent, saturated incidence rate to capture the cumulative effects of infection dynamics and behavioral adaptations in response to an epidemic. This model has previously been successfully employed to analyze the effects of partial lockdown strategies in various European countries. For the purpose of demonstrating the proposed data augmentation strategy, we select a dataset corresponding to a period marked by COVID-19 lockdowns. It is worth to mention that, the complexity introduced by lockdown measures, makes the forecasting process more challenging compared to standard scenarios. Therefore, in the following, we employ system (2.2) to develop our data-driven strategy and the associated deep learning framework, enabling epidemic forecasting.
@@ 3 On the the construction of a data driven social-SIR model
In the context of epidemic prediction, mathematical models such as those presented in the previous section provide valuable insights into the spread of infectious diseases. However, these models often rely on predefined assumptions and parameters that may not fully capture the complexities of real-world epidemics. Neural networks offer a powerful alternative by learning directly from data, enabling them to model complex, nonlinear relationships that traditional epidemic models may struggle to represent. However, one major limitation of training neural networks exclusively on available data is that, while they can effectively interpolate within a given time interval, they struggle to make accurate predictions beyond the observed data points [33, 2, 9]. To address this issue, we propose an approach that combines a mathematical epidemiological model—the social-SIR model (2.2)—with real outbreak data. The key idea is to generate additional synthetic information by solving this generalized SIR model, thereby filling the gaps between available real or measured data points. With this enriched dataset, one can train a neural network more effectively to predict future infection trends in a given population. As we demonstrate later, this approach leverages both real and synthetic data, significantly improving predictive accuracy and enhancing the model’s ability to forecast epidemic dynamics. However, to implement this strategy, the model must be capable of interpolating through the experimental data points. Therefore, in the following section, we discuss how the parameters in (2.2) are defined.
@@ 3.1 Parameters estimation
To develop our approach, we focus on the COVID-19 outbreak in two European countries, Italy and Spain, during the 2020 pandemic. Recognizing that data collection processes inherently introduce errors, particularly due to underreporting of number of infected individuals in the pandemic’s early stages, we assume an intrinsic uncertainty is embedded in the collected data, capturing the variability and limitations of real-world measurements, [18]. While the development of advanced uncertainty quantification techniques falls outside the scope of this work, it represents a promising avenue for future research. In the present study, we instead adopt the well?established framework described in [23, 1]. Given that both countries implemented restrictive measures during this period, we proceed as follows. We define two optimization phases: in the first phase, we estimate the parameters $\beta$ and $\gamma$ for the pre-lockdown period using the model (2.2) with $H(t,I(t))=1$; in the second phase, we estimate the function $H(t,I(t))$ from lockdown and post-lockdown data.
@@
Both the estimated transmission rates $\beta_{est}$ and the recovery rates $\gamma_{est}$ have units of inverse time (e.g., $\texttt{day}^{-1}$), ensuring the correct dimensional consistency of the model equations. In Figure 1 on the left the evolution of the dynamics in the pre-lockdown phase in Italy obtained by solving the social-SIR model (2.2) with $H(t,I(t))=1$, using the estimated parameters $\beta=\beta_{est}^{It}$ and $\gamma=\gamma_{est}^{It}$ in (3.4). This is compared, in the same image, to the experimental data from the pre-lockdown phase, spanning up to time $t_{l}=14$ with a time step $\Delta t=1$ corresponding to $14$ days with $1$ day interval between the measurements. The estimated dynamics aligns quite well with the available data, as confirmed by the embedded plot which shows the pointwise error over time between the solution of the generalized SIR model and the data at disposal, i.e.
@@
Once $\beta$ and $\gamma$ have been estimated, we proceed to determine the pointwise values $H(t,I(t))$ in (2.2) in such a way that the model fits as well as possible the lockdown and post-lockdown data, i.e., data in the interval $[t_{l},T]$, where $T$ (set to July 1st) is the final observation time for both countries. To this end, we solve the following new optimization problem
@@
where $\mathbb{P}_{[0,1]}(\cdot)$ denotes the projection over the interval $[0,1]$. As demonstrate later, the function $H$ computed as in (3.12) tends to exhibit oscillatory behavior over time. To reduce these oscillations, we propose to replace the projection operator $\mathbb{P}_{[0,1]}(\cdot)$ in equation (3.12) with a smoother approximation of the type
$\mathbb{S}(x)=\frac{1}{1+e^{-kx}},$ (3.13)
for a certain $k>0$, and $x\in\mathbb{R}$. This modification helps to regularize the value of $H$. The results of this study are reported in Figure 2 where on the left we show the lockdown and post-lockdown experimental data and the solution of the generalized SIR model with the contact function $H$ as in (3.12) and, alternatively, with the smoothed version (3.13) setting $k=1.6$. The estimated and true solutions are in good agreement as shown in the embedded plot which represents the corresponding pointwise error computed as in (3.6). On the right, we plot the estimated values of $H^{n}$ over time, comparing the values obtained from the original projection-based approach and the smoothed operator. The function $H$ computed as in (3.12) exhibits oscillations over time, suggesting potential measurement errors in the data such as a non-exponential decay of the number of infected especially in the late post-lockdown phase, mimicking the intrinsic uncertainty that may affect data. These oscillations are significantly reduced when using the smoothed operator. However, a reduction in accuracy of the numerical solution can be observed. First row, Italy. Second row, Spain.
Figure 2: Lockdown and post-lockdown phase. On the left the dynamics of the infected population computed by solving the generalized SIR model (2.2) with $\beta$ and $\gamma$ as in (3.4) (Italy), and as in (3.5) (Spain), assuming the contact function $H$ to be defined as in (3.12) and, alternatively, with the smoothed version (3.13) setting $k=1.6$ compared with the experimental data. In the embedded plot the pointwise error between the two infected populations computed as in (3.6). On the right the contact function $H$ computed as in (3.12) and its smoothed version computed introducing an operator as in (3.13). First row, Italy. Second row, Spain.
@@ 4.1 Neural network for epidemic prediction
To evaluate the performances of FFNs and NARs in epidemic prediction, we conduct some numerical experiments using both real and synthetic data. The real data covers the period from the lockdown start date (in Italy March 9th, in Spain March 14th) to July 1st. These data are available on a daily basis, corresponding to a time step of $\Delta t_{r}=1$. The synthetic data are generated by solving the generalized SIR model given in (2.2), with $\beta$ and $\gamma$ as in (3.4) (Italy), and as in (3.5) (Spain) over the same time interval, using a finer time step $\Delta t_{s}=0.2$, and with $H$ estimated by solving the minimization problem in (3.7). In the following, we will focus on the solution generated by solving the social-SIR model with $H$ as in (3.12) to generate synthetic data. The version with the smoothing operator is not considered further due to the loss of accuracy it introduces. Furthermore, since $H$ is originally defined at daily intervals as in (3.12), we use linear interpolation to extend its definition to the finer grid. Both networks are trained using the Levenberg-Marquardt optimizer for $500$ epochs. In the case of real data we set $\omega_{r}=1$ and $\omega_{s}=0$, while in the case of synthetic data we set $\omega_{r}=0$ and $\omega_{s}=1$ in (4.3).
@@
We design a FFN with three fully connected layers, each consisting of $10$ nodes, and use the tanh activation function, relying on the use of the Matlab function feedforwardnet. The input to the network is the time interval $[t_{l},T]$, being $t_{l}$ the lockdown starting date and $T$ the final time supposed to be July 1st. The output is the normalized number of infected $I(t)$, which corresponds to either real data or synthetic data generated as described above. We train the neural network using the Matlab function train. To evaluate the neural network’s performance, we first compute its solution over the training set for both Italy and Spain. Specifically, we consider the time interval $[t_{l},T]$ with a time step $\Delta t_{r}=1$. In the left panel of Figure 3 and of Figure 4, we compare the solution obtained using a neural network trained on both real and synthetic data, with the available data for respectively the Italian and the Spanish case. The neural network provides a good approximation of the solution, making it well-suited for interpolation tasks. Next, to assess its predictive capabilities, we focus on the time interval $(T,T_{test}]$, where $T$ (July 1st) is the starting time and $T_{test}$ (July 15th) is the final time, using the same daily discretization. The right panels of Figures 3 and 4 present a comparison between the network’s predictions for the test set and the daily real data available in that period. Additionally, the figures display the pointwise error between the neural network’s outputs and the real data, calculated as in (3.6). The standard FNN as expected faces challenges in forecasting future dynamics and the data augmentation does not provide a clear improvement in the results.